OpenStudy (anonymous):

Evaluate the function f(x)=7x+1 at the given values of the independent variable and simplify. f(3)= f(x+4)= f(-x)=

OpenStudy (michele_laino):

hint: \[f\left( 3 \right) = \left( {7 \times 3} \right) + 1 = ...?\]

OpenStudy (michele_laino):

I replaced x with 3